OpenStudy (anonymous):

point P(2,5) is on the terminal arm of an angle in standard position a) Determine the distance from the origin P(radical) b)write the primary trigonometric ratios of c) what is the measure of an angle to the nearest degree

OpenStudy (anonymous):

@Psymon can u help me to see if my answers are right

OpenStudy (psymon):

I will shortly, just in another problem :3

OpenStudy (anonymous):

 i draw this triangle to help me

OpenStudy (anonymous):

so the distance is square root of 29 and my sin is 5 \[\sqrt{29}\] and my cos is 2/\[\sqrt{29}\] and my tan 2/5 so my angle is 22 ° i guess

OpenStudy (psymon):

Tangent is 5/2, not 2/5, you reversed it :P So from there we could say that tanx = 5/2, meaning tan^-1(5/2) = x

OpenStudy (anonymous):

ohh yeah so my angles is 68°, again thank you so much

OpenStudy (psymon):

Yep, good job :3

OpenStudy (anonymous):

seriously thank you

OpenStudy (psymon):

Yeah, glad I can help : )

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

can u help me. to solve this one, Determine the slope of the terminal arm for an angle =53° in standard position give the answer to 1 decimal place... please

OpenStudy (psymon):

Yeah, this one takes more logic to figure out. Basically, no matter what the side lengths are, the slope of a 53 degree angle will always be the same. So what Im going to do is just pick the number 10 randomly and Ill draw this picture:  Now slope is rise over run, so up and down the y axis, right left on the x-axis. Because of that, I dont care at all about the hypotenuse. The hypotenuse doesnt give me enough information. So I picked a random number 10 for my "run", and now Im going to solve for the opposite side, my "rise". SO given the drawing, do you know which trig function to use to help me find x?